OpenStudy (anonymous):

when baking a cake,you have the choice of the following pans,a round one that is 2 inches deep and has a 7 inch diameter and 6x9 inch rectangular cake pan that is 2 inches deep. A.which of these pans are bigger b.excluding the bottom of the cake what is the surface area of each cake that will be covered in frosting?justify your answer

OpenStudy (anonymous):

its a tough one

OpenStudy (perl):

by bigger we mean which pan has more volume (space)

OpenStudy (perl):

$$ \Large{ Volume~ of ~{cylinder} = \pi r^2h \\Volume ~ of ~ rectangle = length \cdot width \cdot height } $$